OpenStudy (anonymous):

The area of a square is found by squaring one of its sides, (A = s 2). A certain square has an area of z 2 + 18z + 81. Factor the trinomial to find a side of the square.

OpenStudy (anonymous):

side = z + 27 side = z + 18 side = z + 9

OpenStudy (chosenmatt):

welcome to openstudy

OpenStudy (zpupster):

we need the factors of 81 1 81 1 3 27 81 1 3 9 27 81 these are the factors lets split the z^2 (z (z what factors of 81 above add to get +18?